An electron travel at 2.0x10^7m/s in a plane perpendicular to a uniform 0.010T magnetic field. Describe its path quantitatively. (Mass of an electrons is 9.1x10^-31 kg, the charge in an electron is 1.6x 10^-19C)

1
Expert's answer
2022-04-17T17:06:16-0400

The particles moves along a circle of radius

"R=\frac{mv}{qB}\\
=\frac{9.1*10^{-31}*2.0*10^7}{1.6*10^{-19}*0.010}=0.011\:\rm m"